# Hybrid Meta-learners for Estimating Heterogeneous Treatment Effects

### v1

@@ Abstract
Estimating conditional average treatment effects (CATE) from observational data involves modeling decisions that differ from supervised learning, particularly concerning how to regularize model complexity. Previous approaches can be grouped into two primary “meta-learner” paradigms that impose distinct inductive biases. Indirect meta-learners first fit and regularize separate potential outcome (PO) models and then estimate CATE by taking their difference, whereas direct meta-learners construct and directly regularize estimators for the CATE function itself. Neither approach consistently outperforms the other across all scenarios: indirect learners perform well when the PO functions are simple, while direct learners outperform when the CATE is simpler than individual PO functions. In this paper, we introduce the Hybrid Learner (H-learner), a novel regularization strategy that interpolates between the direct and indirect regularizations depending on the dataset at hand. The H-learner achieves this by learning intermediate functions whose difference closely approximates the CATE without necessarily requiring accurate individual approximations of the POs themselves. We demonstrate empirically that intentionally allowing suboptimal fits to the POs improves the bias-variance tradeoff in estimating CATE. Experiments conducted on semi-synthetic and real-world benchmark datasets illustrate that the H-learner consistently operates at the Pareto frontier, effectively combining the strengths of both direct and indirect meta-learners. Code: https://github.com/AlaaLab/H-learner
@@ 1 Introduction
We consider the problem of estimating conditional average treatment effects (CATE) from observational data [1, 2, 3, 4, 5, 6, 7]. Let $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ denote the observational dataset, where $X_{i}\in\mathcal{X}$ is a feature vector, $T_{i}\in\{0,1\}$ is a binary treatment indicator and $Y_{i}\in\mathbb{R}$ is the observed outcome of interest. Under the Neyman-Rubin potential outcomes framework [8], let $Y_{i}(1)$ and $Y_{i}(0)$ denote the potential outcomes (POs) for individual $i$ under treatment and control, respectively. The fundamental problem of causal inference is that we only observe one of the POs for each individual, i.e., $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. We denote $\mu_{0}(x)=\mathbb{E}[Y(0)\mid X=x]$ and $\mu_{1}(x)=\mathbb{E}[Y(1)\mid X=x]$ as the expected POs under control and treatment. Our goal is to estimate the CATE function:
@@
[…]
@@
Learning the CATE function differs from standard supervised learning in several important ways. As noted by [10], two key distinctions stand out: (1) covariate shift induced by confounding, where the feature distributions differ across treatment groups, i.e., $\mathbb{P}(X\mid T=1)\neq\mathbb{P}(X\mid T=0)$, and (2) the inductive biases associated with the goal of estimating a difference between two outcome functions, $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$, rather than the individual POs, which introduces unique considerations in how we design and regularize the models for $\mu_{0}(x)$, $\mu_{1}(x)$, and $\tau(x)$. In this paper, we focus on the second aspect of inductive bias and propose a new strategy for regularizing CATE models.
To understand how regularization affects the learning of CATE, consider the following illustrative example (variants of this example appear in the literature [4, 5, 10]). We define the POs as:
@@
Here, the control and treatment outcome functions share the same sinusoidal shape and frequency, and differ only by a phase shift $\Delta$ and an amplitude shift $\beta$. As a result, both $\mu_{0}(x)$ and $\mu_{1}(x)$ have comparable complexity, i.e., both are equally easy—or difficult—to learn. However, the complexity of the resulting treatment effect function $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$ can vary significantly depending on the value of $\Delta$. We highlight this by considering the following two scenarios:
@@
This example illustrates that the intrinsic difficulty of learning the treatment effect depends not just on the complexity of the POs, but also on the relationship between them. As such, choosing how to regularize CATE estimators requires careful consideration of which function—outcome or effect—is likely to be simpler in the given domain.
Künzel et al. [5] coined the notion of “meta-learners” to describe a family of model-agnostic strategies for learning CATE. Broadly, these strategies fall into two categories [10]:
@@
Direct meta-learners target the CATE function directly by first constructing a pseudo-outcome $Y_{\varphi}$ based on some nuisance parameter $\varphi$ (e.g., estimates of the propensity score $\widehat{\pi}(x)$ and the POs $\widehat{\mu}_{t}$, $t\in\{0,1\}$), and then fit a model $\widehat{\tau}(x)$ for the CATE function using the dataset $\{(X_{i},Y_{\varphi,i})\}$.
@@
Figure 1 illustrates how direct and indirect meta-learners perform under Scenarios 1 and 2. In Scenario 1, where the CATE function is much simpler than either PO, the direct learner outperforms the indirect approach. This is because indirect learners fit and regularize the two PO models separately, and each model tends to make errors in different regions of the covariate space $\mathcal{X}$. When the two models are subtracted to estimate the treatment effect, they can compound in unpredictable ways, introducing spurious heterogeneity into the implied CATE function. This phenomenon is known in the literature as “regularization-induced confounding” [11, 12]. Conversely, in Scenario 2—where the CATE function is as complex as the POs—direct learners tend to underperform, as the benefit of targeting a simpler effect function no longer holds. In practice, however, it is difficult to know in advance which strategy will perform better, as their relative performance depends on sample size, treatment imbalance, the degree of confounding and overlap, and the specific construction of pseudo-outcomes.
@@
We propose a novel regularization strategy for CATE estimation that generalizes both direct and indirect meta-learners. Our strategy formulates a proximal indirect learning task, where the model learns two intermediate functions, $f_{0}$ and $f_{1}$, which are not necessarily optimal estimators of the POs but are trained such that their difference approximates the treatment effect. Regularization is applied by weighting two competing objectives: encouraging $f_{0}$ and $f_{1}$ to closely approximate the true POs, and ensuring that their difference $f_{1}-f_{0}$ aligns with a pseudo-outcome-based guess of CATE.
Following the taxonomy established in the literature, we refer to our method as the Hybrid learner (H-learner), as it combines elements of both direct and indirect regularization strategies. Figure 2 illustrates the performance of the H-learner under the two scenarios described earlier. In Scenario 1, the H-learner identifies intermediate functions $\widehat{f}_{0}$ and $\widehat{f}_{1}$ that are not perfect fits to the POs, but whose difference is regularized by what a direct estimator would have produced. This leads to a more accurate CATE estimate than those obtained by either direct or indirect learners alone. In Scenario 2, the H-learner adapts to the more complex treatment effect structure by placing greater emphasis on accurately fitting the two POs, and continues to perform competitively, even when compared to an
Figure 2: Depiction of the PO and CATE estimates based on the H-learner regularization strategy.
indirect learner, which focuses solely on outcome prediction.
In the remainder of the paper, we describe the proposed H-learner regularization strategy and place it in the context of existing meta-learning approaches for CATE estimation. Our method is model-agnostic and compatible with a wide range of machine learning architectures. We demonstrate that H-learners improve the bias–variance tradeoff in estimating CATE. Experiments on semi-synthetic and real-world datasets show that it consistently lies on the Pareto frontier by adapting the strengths of direct and indirect learners to different data generation processes (DGPs).
@@ 2 Meta-Learners for Estimating Heterogeneous Treatment Effects
In this section, we describe direct and indirect meta-learners, illustrate each with representative examples from the literature, and discuss the implicit inductive biases and regularization they impose.
@@ 2.1 Indirect Meta-learners
Given an observational dataset partitioned by the two treatment groups, $\mathcal{D}_{0}=\{(X_{i},Y_{i}):T_{i}=0\}$ and $\mathcal{D}_{1}=\{(X_{i},Y_{i}):T_{i}=1\}$, an indirect meta-learner estimates the PO models $\mu_{0}$ and $\mu_{1}$ by independently fitting outcome models on each group by minimizing a regularized empirical risk of the form:
@@
Several well-known methods fall under the category of indirect meta-learners. The T-learner fits two separate models $\mu_{0}(x;\theta_{0})$ and $\mu_{1}(x;\theta_{1})$ independently using the datasets $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, and estimates CATE as the difference between the predicted POs. Another architecture is the S-learner, which fits a single model $\mu(x,t;\theta)$ by treating the treatment indicator as an additional feature dimension, then estimating CATE by toggling this feature $t$ for each new $x$, i.e., $\mu(x,1;\widehat{\theta})-\mu(x,0;\widehat{\theta})$. Examples of these models include Bayesian additive regression trees (BARTs) and regression trees [13, 14]. Finally, there has been a few representation-based methods such as TARNet [2], which learn a shared latent representation of the features before fitting separate heads for each treatment arm.
A common characteristic of indirect meta-learners trained using the loss in (3) is that the potential outcome models $\mu_{0}$ and $\mu_{1}$ are optimized and regularized independently, each to favor a simpler function that best fits the data in $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, respectively. As noted by [11, 12], this independent regularization can implicitly encourage spurious treatment effect heterogeneity, even when the data do not support it. This occurs because the models are not regularized with respect to the difference that defines the CATE, i.e., $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$. [10] further points out that this inductive bias not only risks introducing artefactual heterogeneity but also contradicts how treatment effect heterogeneity should be approached from a scientific standpoint—where the default (null) hypothesis is often that treatment effects are homogeneous [15, 16]. To encode this prior belief, one must explicitly encourage similarity between $\mu_{0}(x)$ and $\mu_{1}(x)$. A simple, though naïve, strategy is to regularize the difference between the two models toward zero:
[…]
However, this strategy can be overly restrictive in settings where treatment effects are non-zero but homogeneous, as it risks suppressing meaningful differences between the POs. Direct meta-learners provide a more flexible alternative by allowing for direct regularization of the CATE function itself.
@@ 2.2 Direct Meta-learners
Direct meta-learners are typically implemented as two-stage procedures. In the first stage, a nuisance parameter $\varphi$—such as the propensity score or outcome regression—is estimated and used to construct a pseudo-outcome $Y_{\varphi}$. In the second stage, a single machine learning model is trained to estimate ${\tau}(x)$ using the pseudo-labeled dataset $\{(X_{i},Y_{\varphi,i})\}$. The pseudo-outcome is designed so that it depends only on observables and satisfies, either exactly or approximately, the following condition:
@@
[…]
where $\ell$ is a loss function (e.g., squared loss) and $\mathcal{R}(\theta)$ is a regularization term applied directly to the CATE function. The general framework described above captures various meta-learners in previous literature,
including the X-learner, the Inverse propensity weighted (IPW) learner and the Doubly-robust (DR) learner (Table 1). When the propensity score is known, the IPW and DR pseudo-outcomes satisfy (5) exactly. The regularization strategy underlying direct
@@
IPW-learner [17] $Y_{\varphi}=\frac{T-\pi(X)}{\pi(X)(1-\pi(X))}Y$
X-learner [5] $Y_{\varphi}=T(Y-\widehat{\mu}_{0}(X))+(1-T)(\widehat{\mu}_{1}(X)-Y)$
DR-learner [4] $Y_{\varphi}=\frac{T-\pi(X)}{\pi(X)(1-\pi(X))}(Y-\widehat{\mu}_{T}(X))+\widehat{\mu}_{1}(X)-\widehat{\mu}_{0}(X)$
@@
meta-learners differs fundamentally from that of indirect learners: rather than independently regularizing the PO models, direct learners apply training and regularization directly to the CATE function. The effectiveness of this approach hinges on multiple factors, including the quality of the pseudo-outcome transformation—i.e., how accurately it reflects the CATE function—and the relative simplicity of the CATE function compared to the underlying POs. When the treatment effect surface is smoother or lower-dimensional, direct learners often produce more efficient and stable estimates. However, poor pseudo-outcome constructions—such as those based on inverse propensity can often suffer from high variance and degrade performance.
@@ 3 Hybrid Meta-learners
The key idea behind the proposed H-learner is to train a pair of models $f_{0},f_{1}\in\mathcal{F}$, where the CATE is estimated as their difference, $f_{1}(x)-f_{0}(x)$, in a manner similar to indirect meta-learners. However, unlike traditional indirect learners, which train $f_{0}$ and $f_{1}$ solely to minimize empirical risk with respect to the POs, the H-learner jointly optimizes these models to balance both accurate prediction of the POs and the accuracy of their implied CATE. As a result, the optimal finite-sample solutions $\widehat{f}_{0},\widehat{f}_{1}\in\mathcal{F}$ may not correspond to the best fits for the POs individually, but may instead favor a pair whose difference yields a more accurate estimate of the implied treatment effect function. Similar to direct meta-learners, the H-learner follows a two-stage procedure outlined below:
@@
Stage 2: Fit two intermediate functions $f_{0},f_{1}\in\mathcal{F}$ by jointly minimizing the empirical risk of each function in predicting the observed (factual) outcomes, along with the empirical risk of their difference in approximating the pseudo-outcome constructed in Stage 1. Formally, the H-learner loss function balances both outcome prediction and treatment effect estimation, and is given by:
$\min_{f_{0},f_{1}}\,\mbox{$\sum_{i=1}^{n}$}(1-\lambda)\cdot\underbrace{\ell\big{(}Y_{i},f_{T_{i}}(X_{i})\big{)}}_{\text{Indirect component $\ell_{I}$}}+\lambda\cdot\underbrace{\left((f_{1}(X_{i})-f_{0}(X_{i}))-Y_{\varphi,i}\right)^{2}}_{\text{Direct component $\ell_{D}$}},\quad\text{where }\ 0\leq\lambda\leq 1$ (8)
and we obtain the final CATE estimator as $\hat{\tau}(x)=\hat{f_{1}}(x)-\hat{f_{0}}(x)$.
The H-learner loss can be viewed as a regularized version of the indirect learning framework, with a regularization term that encourages the difference between the two intermediate models to align with a pseudo-outcome. Alternatively, it can be interpreted as a generalized regularization strategy that interpolates between direct and indirect learners and adaptively balances the influence of each. For example, in cases where the CATE function is significantly simpler than the POs—such as Scenario 1 in Section 1—setting $\lambda=1$ reduces the H-learner to a direct learner, which is optimal in that setting (Figure 3(b)). Conversely, in Scenario 2, where the CATE is as complex as the outcome functions, setting $\lambda=0$ corresponds to a fully indirect learner and may yield better results. In practice, the optimal choice of $\lambda$ typically lies somewhere in between 0 and 1, reflecting the benefit of leveraging both regularization strategies rather than relying on either one alone. This optimal balance $\lambda$ depends on various underlying factors, including the relative complexity of the CATE and PO functions. We treat $\lambda$ as a hyperparameter and adaptively tune its value by minimizing the validation loss. This adaptivity allows H-learner to dynamically adjust its inductive bias to better align with the underlying DGP, resulting in consistently improved performance across a wide range of scenarios.
@@
Input :   $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$
Output :   $\widehat{\tau}(x)=\widehat{f}_{1}(x)-\widehat{f}_{0}(x)$
1
2 Split $\mathcal{D}$ into $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$;
3 Stage 1: Pseudo-outcome construction
5 Construct ${Y}_{\varphi}$ via the options in Table 1;
6 Stage 2: Training the H-learner
Return $\widehat{\tau}(x)=\widehat{f}_{1}(x)-\widehat{f}_{0}(x)$
Algorithm 1 H-learner
The H-learner procedure is compatible with any existing CATE estimation architecture that provides explicit estimates of the POs. In our implementation (Algorithm 1), we use the TARNet architecture both to estimate the nuisance components required for constructing pseudo-outcomes and to fit the intermediate models used to derive the final CATE estimate as illustrated in Figure 3(a). Further implementation details can be found in the supplementary material.

### v2

@@ Abstract
Estimating conditional average treatment effects (CATE) from observational data involves modeling decisions that differ from supervised learning, particularly concerning how to regularize model complexity. Previous approaches can be grouped into two primary “meta-learner” paradigms that impose distinct inductive biases. Indirect meta-learners first fit and regularize separate potential outcome (PO) models and then estimate CATE by taking their difference, whereas direct meta-learners construct and directly regularize estimators for the CATE function itself. Neither approach consistently outperforms the other across all scenarios: indirect learners perform well when the PO functions are simple, while direct learners outperform when the CATE is simpler than individual PO functions. In this paper, we introduce the Hybrid Learner (H-learner), a novel regularization strategy that interpolates between the direct and indirect regularizations depending on the dataset at hand. The H-learner achieves this by learning intermediate functions whose difference closely approximates the CATE without necessarily requiring accurate individual approximations of the POs themselves. We demonstrate that intentionally allowing suboptimal fits to the POs improves the bias-variance tradeoff in estimating CATE. Experiments conducted on semi-synthetic and real-world benchmark datasets illustrate that the H-learner consistently operates at the Pareto frontier, effectively combining the strengths of both direct and indirect meta-learners. Code: https://github.com/AlaaLab/H-learner
@@ 1 Introduction
We consider the problem of estimating conditional average treatment effects (CATE) from observational data (Johansson et al., 2016; Alaa and Van Der Schaar, 2017; Shalit et al., 2017; Wager and Athey, 2018; Künzel et al., 2019; Kennedy, 2020). Let $\mathcal{D}=\{(X_{i},T_{i},Y_{i})\}_{i=1}^{n}$ denote the observational dataset, where $X_{i}\in\mathcal{X}$ is a feature vector, $T_{i}\in\{0,1\}$ is a binary treatment indicator and $Y_{i}\in\mathbb{R}$ is the observed outcome of interest. Under the Neyman-Rubin potential outcomes framework (Neyman, 1923), let $Y_{i}(1)$ and $Y_{i}(0)$ denote the potential outcomes (POs) for individual $i$ under treatment and control, respectively. The fundamental problem of causal inference is that we only observe one of the POs for each individual, i.e., $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. We denote $\mu_{0}(x)=\mathbb{E}[Y(0)\mid X=x]$ and $\mu_{1}(x)=\mathbb{E}[Y(1)\mid X=x]$ as the expected POs under control and treatment. Our goal is to estimate the CATE function:
@@
using the dataset $\mathcal{D}$. This estimand is identifiable from the sample $\mathcal{D}$ under the following assumptions: (1) Consistency: $Y_{i}=Y_{i}(t)$, (2) Unconfoundedness: $(Y(0),Y(1))\perp\!\!\!\perp T\mid X$, and (3) Positivity: $\forall x\in\mathcal{X},\ 0<\pi(x)<1$, where $\pi(x)=\mathbb{P}(T=1\mid X=x)$ is the propensity score (Rubin, 1974).
@@
Learning the CATE function differs from standard supervised learning in several important ways. As noted by Curth and van der Schaar (2021), two key distinctions stand out: (1) covariate shift induced by confounding, where the feature distributions differ across treatment and control groups, i.e., $\mathbb{P}(X\mid T=1)\neq\mathbb{P}(X\mid T=0)$, and (2) the inductive biases associated with the goal of estimating a difference between two outcome functions, $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$, rather than the individual POs, which introduces unique considerations in how we design and regularize the models for $\mu_{0}(x)$, $\mu_{1}(x)$, and $\tau(x)$. In this paper, we focus on the second aspect and propose a new strategy for regularizing CATE models.
To understand how regularization affects learning of CATE, consider the following illustrative example, with variants of this example appear in the literature (Künzel et al., 2019; Kennedy, 2020; Curth and van der Schaar, 2021). We define the POs as:
@@
Here, the control and treatment outcome functions share the same sinusoidal shape and frequency, and differ only by a phase shift $\Delta$ and an amplitude shift $\beta$. As a result, both $\mu_{0}(x)$ and $\mu_{1}(x)$ have comparable complexity, i.e., both are equally easy—or difficult—to learn. However, the complexity of the resulting treatment effect function $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$ can vary significantly depending on the values of $\Delta$. We highlight this by considering the following two scenarios:
@@
This example illustrates that the intrinsic difficulty in learning the treatment effect depends not just on the complexity of the POs, but also on the relationship between them. As such, choosing how to regularize CATE estimators requires careful consideration of which function—outcome or effect—can be better estimated in the given domain.
Künzel et al. (2019) coined the notion of “meta-learners” to describe a family of model-agnostic strategies for learning CATE. Broadly, these strategies fall into two categories (Curth and van der Schaar, 2021):
@@
Direct meta-learners target the CATE function directly by constructing a pseudo-outcome $Y_{\varphi}$ based on some nuisance parameter $\varphi$ (e.g., estimates of the propensity score $\widehat{\pi}(x)$ and the POs $\widehat{\mu}_{t}$, $t\in\{0,1\}$), and then fit a model $\widehat{\tau}(x)$ for the CATE function using the constructed dataset $\{(X_{i},Y_{\varphi,i})\}$.
@@
Figure 1 shows how direct and indirect learners perform under both scenarios. In Scenario 1, where the CATE function is simpler than the PO, the direct learner outperforms the indirect approach. This is because indirect learners separately fit and regularize the PO models, and each model make errors in different regions of the covariate space. When the two models are subtracted to estimate the treatment effect, they can compound in unpredictable ways, introducing spurious heterogeneity into the CATE function. This phenomenon is known as “regularization-induced confounding” (Hahn et al., 2018). Conversely, in Scenario 2, where the CATE function is as complex as the POs, direct learners tend to underperform, as the benefit of targeting a simpler effect function no longer holds. In practice, however, it is difficult to know in advance which strategy will perform better, as their relative performance also depends on other factors underlying data generation processes (DGPs), including sample sizes, treatment imbalance, the degree of confounding and overlap.
@@
To address these limitations, we propose the Hybrid learner (H-learner), a novel meta-learner regularization strategy for CATE estimation that generalizes both direct and indirect meta-learners. Our strategy formulates a proximal indirect learning task, where the model learns two intermediate functions, $f_{0}$ and $f_{1}$, which are not necessarily optimal estimators of the POs but are trained such that their difference accurately approximates the treatment effect. Regularization is applied by weighting two competing objectives: encouraging $f_{0}$ and $f_{1}$ to closely approximate the true POs, and ensuring that their difference $f_{1}-f_{0}$ aligns with a pseudo-outcome-based guess of CATE.
In the remainder of the paper, we describe the proposed H-learner regularization strategy and place it in the context of existing meta-learning paradigms. Our method is model-agnostic and compatible with a wide range of machine learning (ML) architectures. In particular, we make the following contributions:
  1. 1.
    We characterize the performance of meta-learners by identifying key factors of the underlying DGPs and explicitly linking them to the inductive biases introduced by their regularization strategies.
  2. 2.
    We introduce the H-learner as a unified approach that combines direct and indirect regularization and theoretically demonstrate that it achieves lower prediction risk by balancing the bias–variance trade-off inherent in existing meta-learners.
  3. 3.
    We empirically show that the H-learner consistently achieves state-of-the-art performance, remains robust across diverse DGP characteristics, and lies on the Pareto frontier of common benchmarks.
@@ 2 Meta-Learners for Estimating Heterogeneous Treatment Effects
In this section, we illustrate representative examples of indirect and direct meta-learners from the literature, and discuss the implicit inductive biases and regularization they impose.
@@ 2.1 Indirect Meta-learners
Given an observational dataset partitioned by the two treatment groups, $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$, an indirect meta-learner estimates the PO models $\mu_{0}$ and $\mu_{1}$ by independently fitting outcome models on each group by minimizing a regularized empirical risk of the form:
$\mbox{$\sum_{i=1}^{n}$}\ell\big(Y_{i},\mu_{T_{i}}(X_{i};\theta_{T_{i}})\big)+\lambda\,\mathcal{R}(\theta_{0},\theta_{1}),$ (3)
@@
However, because indirect learners optimize and regularize $\mu_{0}$ and $\mu_{1}$ independently, their inductive bias favors simpler functions that best fit $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$. As a result, taking their difference can yield biased estimates and implicitly induce spurious treatment effect heterogeneity, particularly under DGPs where the CATE function is simpler than the POs (Hahn et al., 2018, 2020). This occurs because the models are not regularized with respect to the difference that defines the CATE, i.e., $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$. Curth and van der Schaar (2021) further points out that this inductive bias not only risks introducing artefactual heterogeneity but also contradicts how treatment effect heterogeneity should be approached from a scientific standpoint—where the default (null) hypothesis is often that treatment effects are homogeneous (Crump et al., 2008; Ballman, 2015).
@@ 2.2 Direct Meta-learners
Direct meta-learners are typically implemented as two-stage procedures. In the first stage, a nuisance parameter $\varphi$—such as the propensity score or outcome regression—is estimated and used to construct a pseudo-outcome $Y_{\varphi}$. In the second stage, a single model is trained to estimate ${\tau}(x)$ using the pseudo-labeled dataset $\{(X_{i},Y_{\varphi,i})\}$. The pseudo-outcome is designed so that it depends only on observables and satisfies, either exactly or approximately, the following condition:
@@
$\mbox{$\sum_{i=1}^{n}$}\ell\big(Y_{\varphi,i},\tau(X_{i};\theta)\big)+\lambda\,\mathcal{R}(\theta),$ (5)
where $\ell$ is a loss function and $\mathcal{R}(\theta)$ is a regularization applied directly to the CATE function. The general framework described above captures various methods in literature, including the X-learner, the Inverse propensity weighted (IPW) learner and the Doubly-robust (DR) learner (Table 1). When the propensity score is known, the IPW and DR pseudo-outcomes satisfy (4) exactly. The regularization strategy underlying direct meta-learners differs fundamentally from that of indirect learners: rather than independently regularizing the PO models, direct learners apply training and regularization directly to the CATE function.
However, the effectiveness of this approach hinges on the quality of the pseudo-outcome transformation, i.e., how accurately it reflects the CATE function, and it is largely determined by factors underlying the DGP. When the CATE surface is higher-dimensional than the underlying POs, direct learners often yield poorer estimates as they target a harder function. Imbalanced propensities can further worsen pseudo-outcome constructions—such as those based on inverse propensity—which suffer from high variance and degrade performance (Curth and Van der Schaar, 2021).
@@
IPW-learner (Horvitz and Thompson, 1952) $Y_{\varphi}=\frac{T-\pi(X)}{\pi(X)(1-\pi(X))}Y$
X-learner (Künzel et al., 2019) $Y_{\varphi}=T(Y-\widehat{\mu}_{0}(X))+(1-T)(\widehat{\mu}_{1}(X)-Y)$
DR-learner (Kennedy, 2020) $Y_{\varphi}=\frac{T-\pi(X)}{\pi(X)(1-\pi(X))}(Y-\widehat{\mu}_{T}(X))+\widehat{\mu}_{1}(X)-\widehat{\mu}_{0}(X)$
@@ 3 Hybrid Meta-learners
The key idea behind the proposed H-learner is to train a pair of models $f_{0},f_{1}\in\mathcal{F}$, where the CATE is estimated as their difference, $f_{1}(x)-f_{0}(x)$. However, unlike traditional indirect learners, which train $f_{0}$ and $f_{1}$ solely to minimize empirical risk with respect to the POs, the H-learner jointly optimizes these models to balance both accurate prediction of the POs and the accuracy of their implied CATE. As a result, the optimal finite-sample solutions $\widehat{f}_{0},\widehat{f}_{1}\in\mathcal{F}$ may not correspond to the best fits for the POs individually, but may instead favor a pair whose difference yields a more accurate estimate of the implied treatment effect function. Similar to direct meta-learners, the H-learner follows a two-stage procedure outlined below (Figure 2):
@@
Stage 2: Fit two intermediate functions $f_{0},f_{1}\in\mathcal{F}$ by jointly minimizing the empirical risk of each function in predicting the observed (factual) outcomes, along with the empirical risk of their difference in approximating the pseudo-outcome constructed in Stage 1. Formally, the H-learner loss function balances both POs prediction and treatment effect estimation, and is given by:
[…]
$+\lambda\,\underbrace{\big((f_{1}(X_{i})-f_{0}(X_{i}))-Y_{\varphi,i}\big)^{2}}_{\text{Direct component $\ell_{D}$}},$
where $0\leq\lambda\leq 1$, and we obtain the final CATE estimator as $\hat{\tau}(x)=\hat{f_{1}}(x)-\hat{f_{0}}(x)$.
The H-learner loss can be viewed as a regularized version of the indirect learning framework, with a regularization term that encourages the difference between the two intermediate models to align with a pseudo-outcome. Alternatively, it can be interpreted as a generalized regularization strategy that interpolates between direct learners ($\lambda=1$) and indirect learners ($\lambda=0$) while adaptively balancing the influence of each. In the following sections, we demonstrate that the optimal value of $\lambda$ often depends on several factors of the underlying DGP, but in most cases, lies strictly within $\lambda\in(0,1)$ rather than at the boundary values. This reflects the benefit of leveraging both regularization strategies rather than relying on either one alone. Practically, we treat $\lambda$ as a hyperparameter and adaptively tune its value by minimizing the validation loss. Details of this validation procedure are provided in Appendix A.
@@
The H-learner procedure is compatible with any existing CATE estimation architecture that provides explicit estimates of the POs. In our implementation, we use the TARNet architecture both to estimate the nuisance components required for constructing pseudo-outcomes and to fit the intermediate models used to derive the final CATE estimate as illustrated in Figure 2(a). Further implementation details with a pseudo-algorithm can be found in the Appendix A.